# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v1

@@ Abstract
If $(M^{n},g)$ is a Riemannian manifold satisfying such curvature bounds for $k=2$, then we show that $M$ is contained in a neighbourhood of controlled width of an isometrically embedded $1$-dimensional sub-manifold. From this, we deduce several metric and topological consequences: $M$ has at most linear volume growth and at most two ends, the first Betti number of $M$ is bounded above by $1$, and there is precise information on elements of infinite order in $\pi_{1}(M)$.
@@
If $k=n={\rm dim}(M)$, so that the integral lower bound is on the scalar curvature, assuming in addition that the $n-2$-Ricci curvature is asymptotically non-negative, then we prove that the dimension drop at large scales improves to $n-2$.
@@ 1 Introduction
Broadly speaking, we show that a lower Ricci curvature bound combined with a positive lower bound on the $k^{th}$-scalar curvature forces the manifold to have $(k-1)$-dimensional behavior at large scales. Moreover, if $k$ is the dimension of the manifold (so that the lower bound is on the scalar curvature), the dimensional bound on the geometry at large scales can be improved to $(k-2)$, matching in this way the dimension-drop appearing in the aforementioned conjectures by Gromov. Our approach relies on techniques from metric geometry in which a limit space arises from a sequence of manifolds (see the foundational papers [ChCo1, Colding1, ChcoStructure2, ChCo3]).
The first main result of the paper is Theorem 1.1, which shows the continuity of certain integral functionals of $\mathsf{R}_{k}$ under Gromov-Hausdorff convergence. We recall that, if Yau conjecture holds, given a sequence $(M_{i},g_{i},p_{i})\to(X,\mathsf{d},x)$, one can define a scalar curvature measure $\mu$ on $X$ as a weak limit $\mathsf{R}_{M_{i}}\mathsf{Vol}_{i}\to\mu$ (see [Naberconjectures, Conjecture 2.18]). Hence, continuity results for $\int\mathsf{R}$ under GH convergence would ensure that, if $X$ is itself a smooth manifold, then $\mu=\mathsf{R}_{X}\mathsf{Vol}$. Theorem 1.1 proves a weaker result along these lines when $X=\mathbb{R}^{n}$ (see Remark 3.5). Other continuity results for the integral of the scalar curvature were previously obtained in the presence of a lower bound on the sectional curvature in [PetruninPoli, Theorem B3] (after Perelman) and [lebedeva2022curvaturetensorsmoothablealexandrov]. We refer to Definition 2.5 for the definition of $(k,\delta)$-symmetric and $\delta$-regular balls in a manifold. Given $a,b\in\mathbb{R}$, we denote $a\wedge b:=\min\{a,b\}$ and $a\vee b:=\max\{a,b\}$.
@@
Theorem 1.2 has far reaching implications both for the metric and the topological properties of manifolds satisfying (i), (ii), or (iii). These are collected in Theorem 1.3 below. We denote by $\mathrm{b}_{1}(M)$ the first Betti number of $M$. Given a loop $\gamma$ in $M$, we denote by $[\gamma]$ its equivalence class in the fundamental group $\pi_{1}(M)$.
@@
We highlight that an analog of Theorem 1.2 is not to be expected in the higher dimensional case (see Remark 5.9) and that the result fails without the lower bounds on the Ricci curvature (see Remark 5.7). The proof of Theorem 1.2 relies on the study of a new geometric condition on length minimizing geodesics in a metric space (see Section 4).
@@
We now consider the topological implications of almost non-negative Ricci curvature and positive $k^{th}$-scalar curvature. This is the content of Theorems 1.5 and 1.6 (compare with [zhu2024twodimensionvanishingsplittingpositive, Theorems 1.3 and 1.7].
@@ Acknowledgments
Part of this research was carried out at the Hausdorff Institute of Mathematics in Bonn, during the trimester program “Metric Analysis”. The authors wish to express their appreciation to the institution for the stimulating atmosphere, and they ackonowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.
@@ 2 Preliminaries
We start by recalling some basic notation and terminology about metric spaces. Throughout the section, $(X,\mathsf{d})$ denotes a proper, complete, separable, length metric space. A metric measure spaces $(X,\mathsf{d},\mathfrak{m})$ is a triple, where $(X,\mathsf{d})$ is a metric space and $\mathfrak{m}$ is a non-negative Borel measure on $X$ which is finite on bounded sets and whose support is the whole $X$. We denote by $\mathsf{H}^{k}$ the $k$-dimensional Hausdorff measure induced by $\mathsf{d}$ on $X$. Given a smooth manifold $(M,g)$ we denote by $\mathsf{d}$ and $\mathsf{Vol}$ respectively the Riemannian distance and volume measuure induced by $g$. Given an open set $\Omega\subset X$, we denote by $\mathsf{Lip}(\Omega)$, $\mathsf{Lip}_{\rm loc}(\Omega)$ and $\mathsf{Lip}_{c}(\Omega)$ respectively the set of Lipschitz, locally Lipschitz and compactly supported Lipschitz functions in $\Omega$. We denote the slope of $f\in\mathsf{Lip}_{\rm loc}(\Omega)$ at $x\in\Omega$ by
@@ 3.2 First geometric applications in almost non-negative Ricci and integrally-positive scalar curvature
We now obtain the first geometric applications of the results of the previous section. We introduce some notation first. After [ShenKricci], we consider bounds on the $k^{th}$ Ricci curvature.
@@
4 The segment neighbourhood property (SNP) and some metric lemmas
The goal of this section is to show that metric spaces satisfying the segment neighbourhood property (see Definition 4.2), are contained in a neighbourhood of controlled width of an isometrically embedded $1$-dimensional manifold (see Theorem 4.17). This will be a key step to prove Theorem 1.2. We mention that a similar (but unrelated) property for metric spaces was studied in [TrianglesKapo].
Throughout the section, all metric spaces $(X,\mathsf{d})$ will be assumed to be proper, complete, separable, and geodesic. If not otherwise specified, all curves are assumed to be of finite length and parametrized by arc-length. The length functional is denoted by $\mathrm{L}(\cdot)$.
@@ 4 The segment neighbourhood property (SNP) and some metric lemmas
In this section we deal with metric spaces satisfying the following condition, called segment neighbourhood property.
Let us first fix some notation. Given a closed set $K\subset X$ and a point $x\in X$, we call the projection of $x$ into $K$ the subset $\pi_{K}(x)\subset K$ defined by
@@
Definition 4.2 (Segment neighbourhood property, (SNP)).
We say that the metric space $(X,\mathsf{d})$ satisfies the segment neighbourhood property (SNP, for short), if there exist $R,D>0$, with $R\geq 20D$, such that the following holds. For every segment $r$ of length greater than $2R$ in $X$ and for every $t\in(R,\mathrm{L}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$.
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP) (see Definition 4.2) and let $r$ be a ray in $X$. If a point $x\in X$ has a projection on $r$ which lies in $(r(R),r(\infty))$, then $\mathsf{d}(x,r)\leq D$. This fact follows applying (SNP) to the segments $[r(0),r(T)]$ for $T>2R$.
Remark 4.4.
If $(X,\mathsf{d})$ has (SNP) for $R,D>0$, then it also satisfies (SNP) for $R^{\prime},D^{\prime}>0$ for every $R^{\prime}\geq R$ and $D^{\prime}\leq D$.
@@
The next lemma contains the key technical properties of metric spaces sarisfying the (SNP) (see Definition 4.2).
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP) (see Definition 4.2). Let $\alpha:[0,\mathrm{L}(\alpha)]\to X$ be a unit speed curve.
@@
We claim that $\tilde{s}<\mathrm{L}(\alpha)$. If this is not the case, then, combining with (31), $\alpha(\mathrm{L}(\alpha))$ admits a projection on $l$ contained in $[l(x_{1}+3D),l(x_{2}-3D)]$. By the (SNP), $\alpha(\mathrm{L}(\alpha))$ is $D$-close to $[l(x_{1}+3D),l(x_{2}-3D)]$, so that it cannot have a projection on $l$ given by $l(u)$ for $u\in[0,R]\cup[\mathrm{L}(l)-R,\mathrm{L}(l)]$, a contradiction. This shows that $\tilde{s}<\mathrm{L}(\alpha)$.
Hence, for every $\epsilon>0$, there is a projection $p_{\epsilon}$ of $\alpha(\tilde{s}+\epsilon)$ on $l$. Let $p$ be a projection of $\alpha(\tilde{s})$ on $l$ lying in $[l(x_{1}),l(x_{2})]$. By the (SNP), it holds $\mathsf{d}(p,\alpha(\tilde{s}))<D$. It follows that $\mathsf{d}(p_{\epsilon},\alpha(\tilde{s}+\epsilon))<D+\epsilon$. Combining these facts, it holds
@@
We now consider item 3. Since the image of $\alpha$ is compact, $\pi_{l}(\alpha)$ is a bounded subset of $l$. Hence, $\pi_{l}(\alpha)\subset[r(0),r(T)]$ for some $T>0$. We consider now the segment $l:=[r(0),r(T+2R]$. It is easy to check that $\pi_{l}(\alpha)=\pi_{r}(\alpha)$. By item 1, $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(R),l(t)]$ or $[l(t),l(\mathrm{L}(l)-R)]$. This last case cannot happen as $[l(t),l(\mathrm{L}(l)-R)]=[r(t),r(T+R)]$ and $\pi_{l}(\alpha)\subset[r(0),r(T)]$. Hence, $\pi_{l}(\alpha)$ (which again coincides with $\pi_{r}(\alpha)$) contains a $3D$-net of $[l(R),l(t)]=[r(R),r(t)]$. ∎
The next few lemmas are needed to prove Proposition 4.11, which gives the macroscopic description of non-compact spaces having the (SNP).
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP) (see Definition 4.2). Let $r$ be a ray in $X$ and let $y\in X$ with $\mathsf{d}(r,y)>D$. Let $t>0$ and let $\gamma$ be a curve connecting $y$ and $r(t)$. Then, for every $s\in[0,t]$, it holds $\mathsf{d}(r(s),\gamma)\leq 2R$.
@@
We parametrize $\gamma$ by arc length in such a way that $\gamma(0)=r(t)$ and $\gamma(\mathrm{L}(\gamma))=y$. By the (SNP) and the fact that $\mathsf{d}(y,r)>D$, it follows that $\pi_{r}(\gamma(\mathrm{L}(\gamma)))\in[r(0),r(R)]$. By Lemma 4.6, $\pi_{r}(\gamma)$ contains a $3D$-net of $[r(R),r(t)]$. Using again the (SNP), it follows that $[r(R),r(t)]$ is in a $7D$-neighbourhood of $\gamma$. The statement then follows. ∎
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP) (see Definition 4.2). If there are two divergent rays in $X$, then $X$ contains a line.
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP) (see Definition 4.2). Let $r$ be a ray in $X$. Let $y\in X$ be a point such that $\mathsf{d}(y,r)>D$. Let $r_{y}$ be a ray obtained as a limit of the segments connecting $y$ and $r(t_{j})$, for some sequence $t_{j}\to+\infty$. Then, $r$ is contained in the $2R$-neighbourhood of $r_{y}$.
@@
The next proposition gives the macroscopic description of non-compact spaces satisfying the (SNP).
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP) (see Definition 4.2). Assume that $X$ is non-compact and it does not contain a line. There exists a ray $r$ in $X$ whose $10R$-neighbourhood contains $X$.
@@
We now prove the statement of the lemma. If $X$ is contained in the $10R$-neighbourhood of $X$ there is nothing to prove. Otherwise, we claim that there exists an element $y\in X$ at maximal distance from $r$. Let $y_{i}\in X$ be a sequence maximizing the distance from $r$. We can assume that $9R\leq\mathsf{d}(y_{i},r)\leq D^{\prime}(X)$. By the (SNP), all projections of $y_{i}$ on $r$ are contained in $[r(0),r(R)]$. It follows that the sequence $\{y_{i}\}_{i\in\mathbb{N}}$ is precompact, so that there exists a limit point $y\in X$ at maximal distance from $r$.
@@
We now turn our attention to the study of compact spaces having the (SNP). The key result in this sense is Proposition 4.16. We first need a few lemmas.
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP). Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $\bar{l}$ be a segment starting from $l(0)$ with $\pi_{l}(\bar{l}(\mathrm{L}(\bar{l})))\subset[l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$. If $\bar{l}$ has a projection on $l$ lying in $[l(3R),l(\mathrm{L}(l)-3R)]$, then $\pi_{l}(\bar{l})$ contains a $3D$-net of $[l(R),l(\mathrm{L}(l)-R)]$.
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP). Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $p\in X$ be a point with $\mathsf{d}(p,l)\geq 10R$ and let $\bar{l}$ be a segment connecting $p$ to $l(0)$. If $\bar{l}$ has a projection on $l$ lying in $[l(3R),l(\mathrm{L}(l)-3R)]$, then $\mathrm{L}(\bar{l})>\mathrm{L}(l)$.
@@
Since $\mathsf{d}(p,l)\geq 10R$, the (SNP) implies that $\pi_{l}(p)\subset[l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$. By Lemma 4.12, there is a point $\bar{l}(s_{0})$ which is $7D$-close to $l(\mathrm{L}(l)-R)$. Using triangle inequality, this implies
@@
Let $(X,\mathsf{d})$ be a compact metric space satisfying the (SNP). Let $l$ be a segment of maximal length in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $p\in X$ be a point with $\mathsf{d}(p,l)\geq 10R$ and let $\gamma_{1},\gamma_{2}$ be segments connecting $p$ respectively to $l(\mathrm{L}(l))$ and $l(0)$. Let $b_{1},b_{2}\in\{\gamma_{1},\gamma_{2},l\}$ with $b_{1}\neq b_{2}$. Then $\pi_{b_{1}}(b_{2})\subset[b_{1}(0),b_{1}(3R)]\cup[b_{1}(\mathrm{L}(b_{1})-3R),b_{1}(\mathrm{L}(b_{1})]$.
@@
We now consider the case $b_{1}\in\{\gamma_{1},\gamma_{2}\}$ and $b_{2}=l$. If the statement fails, $l$ has a projection on $b_{1}$ lying in $(b_{1}(3R),b_{1}(\mathrm{L}(b_{1})-3R))$. Consider the orientation of $l$ such that $l(0)\in b_{1}$. Then, $\pi_{b_{1}}(l(\mathrm{L}(l)))\subset[b_{1}(0),b_{1}(R)]\cup[b_{1}(\mathrm{L}(b_{1})-R),b_{1}(\mathrm{L}(b_{1})]$ since otherwise, using the (SNP), it follows that $l$ is shorter than $b_{1}$. Hence, by Lemma 4.12, there exists $s>0$ such that $l(s)$ is $R+7D$-close to $p$. This contradicts that $\mathsf{d}(p,l)\geq 10R$.

### v2

@@ Abstract
If $(M^{n},g)$ is a Riemannian manifold satisfying such curvature bounds for $k=2$, then we show that $M$ is contained in a neighbourhood of controlled width of an isometrically embedded $1$-dimensional sub-manifold. From this, we deduce several metric and topological consequences: $M$ has at most linear volume growth and at most two ends, it has bounded 1-Urysohn width, the first Betti number of $M$ is bounded above by $1$, and there is precise information on elements of infinite order in $\pi_{1}(M)$.
@@
If $k=n={\rm dim}(M)$, so that the integral lower bound is on the scalar curvature, assuming in addition that the $(n-2)$-Ricci curvature is asymptotically non-negative, then we prove that the dimension drop at large scales improves to $n-2$.
@@ 1 Introduction
Broadly speaking, we show that a lower Ricci curvature bound combined with a positive lower bound on the $k^{th}$-scalar curvature forces the manifold to have $(k-1)$-dimensional behavior at large scales. Moreover, if $k$ coincides with the dimension of the manifold (so that the lower bound is on the scalar curvature), the dimensional bound on the geometry at large scales can be improved to $(k-2)$, matching in this way the dimension-drop appearing in the aforementioned conjectures by Gromov. Our approach relies on techniques from metric geometry in which a limit space arises from a sequence of manifolds (see the foundational papers [ChCo1, Colding1, ChcoStructure2, ChCo3]).
The first main result of the paper is Theorem 1.1, which shows the continuity of certain integral functionals of $\mathsf{R}_{k}$ under Gromov-Hausdorff convergence. We recall that, if Yau conjecture holds, given a sequence $(M_{i},g_{i},p_{i})\to(X,\mathsf{d},x)$, one can define a scalar curvature measure $\mu$ on $X$ as a weak limit $\mathsf{R}_{M_{i}}\mathsf{Vol}_{i}\to\mu$ (see [Naberconjectures, Conjecture 2.18]). Hence, continuity results for $\int\mathsf{R}$ under GH convergence would ensure that, if $X$ is itself a smooth manifold, then $\mu=\mathsf{R}_{X}\mathsf{Vol}$. Theorem 1.1 proves a result along these lines when $X=\mathbb{R}^{n}$ (see Remark 3.5). Other continuity results for the integral of the scalar curvature were previously obtained in the presence of a lower bound on the sectional curvature in [PetruninPoli, Theorem B3] (after Perelman) and [lebedeva2022curvaturetensorsmoothablealexandrov]. We refer to Definition 2.5 for the definition of $(k,\delta)$-symmetric and $\delta$-regular balls in a manifold. Given $a,b\in\mathbb{R}$, we denote $a\wedge b:=\min\{a,b\}$ and $a\vee b:=\max\{a,b\}$.
@@
Theorem 1.2 has far reaching implications both for the metric and the topological properties of manifolds satisfying (i), (ii), or (iii). These are collected in Theorem 1.3 below. We denote by $\mathrm{b}_{1}(M)$ the first Betti number of $M$. Given a loop $\gamma$ in $M$, we denote by $[\gamma]$ its equivalence class in the fundamental group $\pi_{1}(M)$. We recall that a metric space $(X,\mathsf{d})$ has $1$-Urysohn width $\leq d$ if there exists a $1$-simplex $Y$, and a continuous map $\pi:X\to Y$, such that $\mathrm{diam}(f^{-1}(y))\leq d$ for every $y\in Y$.
@@
We highlight that an analog of Theorem 1.2 is not to be expected in the higher dimensional case (see Remark 5.9) and that the result fails without the lower bounds on the Ricci curvature (see Remark 5.7). The proof of Theorem 1.2 relies on the study of a geometric condition on length minimizing geodesics in a metric space that was first studied in [TrianglesKapo] (see Section 4).
@@
We now consider the topological implications of almost non-negative Ricci curvature and positive $k^{th}$-scalar curvature. This is the content of Theorems 1.5 and 1.6 (compare with [zhu2024twodimensionvanishingsplittingpositive, Theorems 1.3 and 1.7], after [PanBlowDown]).
@@ Acknowledgments
Part of this research was carried out at the Hausdorff Institute of Mathematics in Bonn, during the trimester program “Metric Analysis”. The authors wish to express their appreciation to the institution for the stimulating atmosphere, and they acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.
@@ 2 Preliminaries
We start by recalling some basic notation and terminology about metric spaces. Throughout the section, $(X,\mathsf{d})$ denotes a proper, complete, separable, length metric space. A metric measure spaces $(X,\mathsf{d},\mathfrak{m})$ is a triple, where $(X,\mathsf{d})$ is a metric space and $\mathfrak{m}$ is a non-negative Borel measure on $X$ which is finite on bounded sets and whose support is the whole $X$. We denote by $\mathsf{H}^{k}$ the $k$-dimensional Hausdorff measure induced by $\mathsf{d}$ on $X$. Given a smooth manifold $(M,g)$ we denote by $\mathsf{d}$ and $\mathsf{Vol}$ respectively the Riemannian distance and volume measure induced by $g$. Given an open set $\Omega\subset X$, we denote by $\mathsf{Lip}(\Omega)$, $\mathsf{Lip}_{\rm loc}(\Omega)$ and $\mathsf{Lip}_{c}(\Omega)$ respectively the set of Lipschitz, locally Lipschitz and compactly supported Lipschitz functions in $\Omega$. We denote the slope of $f\in\mathsf{Lip}_{\rm loc}(\Omega)$ at $x\in\Omega$ by
@@ 3.2 First geometric applications in almost non-negative Ricci and integrally-positive scalar curvature
We now obtain the first geometric applications of the results of the previous section. We introduce some notation first (cf. [ShenKricci]).
@@
4 Thin metric spaces
The goal of this section is to show that thin metric spaces (see Definition 4.2, after [TrianglesKapo]) are contained in a neighbourhood of controlled width of an isometrically embedded $1$-dimensional manifold (see Theorem 4.16). This will be a key step to prove Theorem 1.2.
In [TrianglesKapo, Section 5], it is shown that a non-compact thin metric space is contained in a finite neighourhood of a ray or a line. Comparing with [TrianglesKapo, Section 5], Theorem 4.16 treats the compact case as well, and provides explicit uniform bounds on the width of the aforementioned neighbourhood. These bounds are needed for the applications in the subsequent sections and, in particular, to prove Theorem 1.2. On the other hand, unlike [TrianglesKapo, Section 5], we only consider metric spaces $(X,\mathsf{d})$ that are proper, complete, separable, and geodesic.
If not otherwise specified, all curves are assumed to be of finite length and parametrized by arc-length. The length functional is denoted by $\mathrm{L}(\cdot)$.
@@ 4 Thin metric spaces
Given a closed set $K\subset X$ and a point $x\in X$, we call the projection of $x$ into $K$ the subset $\pi_{K}(x)\subset K$ defined by
@@
Definition 4.2 (Thin metric spaces).
Let $R,D>0$ with $R\geq 20D$. We say that the metric space $(X,\mathsf{d})$ is $(R,D)$-thin if for every segment $r$ of length greater than $2R$ in $X$, and for every $t\in(R,\mathrm{L}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$.
In [TrianglesKapo], a metric space is defined to be $R$-thin if it is $(R,R)$-thin according to Definition 4.2. We consider the extra parameter $D>0$ and we restrict ourselves to the case $R\geq 20D$ as this will be sufficient for the subsequent applications.
@@
Let $(X,\mathsf{d})$ be $(R,D)$-thin (see Definition 4.2) and let $r$ be a ray in $X$. If a point $x\in X$ has a projection on $r$ which lies in $(r(R),r(\infty))$, then $\mathsf{d}(x,r)\leq D$. This fact follows by applying Definition 4.2 to the segments $[r(0),r(T)]$ for $T>2R$.
The next lemma is needed to prove Lemma 4.5, which is the key technical tool to study $(R,D)$-thin metric spaces.
@@
The next lemma contains the key technical properties of $(R,D)$-thin metric spaces (see Definition 4.2).
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space (see Definition 4.2). Let $\alpha:[0,\mathrm{L}(\alpha)]\to X$ be a unit speed curve.
@@
We claim that $\tilde{s}<\mathrm{L}(\alpha)$. If this is not the case, then, combining with (31), $\alpha(\mathrm{L}(\alpha))$ admits a projection on $l$ contained in $[l(x_{1}+3D),l(x_{2}-3D)]$. Since $X$ is $(R,D)$-thin, $\alpha(\mathrm{L}(\alpha))$ is $D$-close to $[l(x_{1}+3D),l(x_{2}-3D)]$, so that it cannot have a projection on $l$ given by $l(u)$ for $u\in[0,R]\cup[\mathrm{L}(l)-R,\mathrm{L}(l)]$, a contradiction. This shows that $\tilde{s}<\mathrm{L}(\alpha)$.
Hence, for every $\epsilon>0$, there is a projection $p_{\epsilon}$ of $\alpha(\tilde{s}+\epsilon)$ on $l$. Let $p$ be a projection of $\alpha(\tilde{s})$ on $l$ lying in $[l(x_{1}),l(x_{2})]$. Since $X$ is $(R,D)$-thin, it holds $\mathsf{d}(p,\alpha(\tilde{s}))<D$. It follows that $\mathsf{d}(p_{\epsilon},\alpha(\tilde{s}+\epsilon))<D+\epsilon$. Combining these facts, it holds
@@
We now consider item 3. Since the image of $\alpha$ is compact, $\pi_{l}(\alpha)$ is a bounded subset of $l$. Hence, $\pi_{l}(\alpha)\subset[r(0),r(T)]$ for some $T>0$. We consider now the segment $l:=[r(0),r(T+2R)]$. It is easy to check that $\pi_{l}(\alpha)=\pi_{r}(\alpha)$. By item 1, $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(R),l(t)]$ or $[l(t),l(\mathrm{L}(l)-R)]$. This last case cannot happen as $[l(t),l(\mathrm{L}(l)-R)]=[r(t),r(T+R)]$ and $\pi_{l}(\alpha)\subset[r(0),r(T)]$. Hence, $\pi_{l}(\alpha)$ (which again coincides with $\pi_{r}(\alpha)$) contains a $3D$-net of $[l(R),l(t)]=[r(R),r(t)]$. ∎
The next few lemmas are needed to prove Proposition 4.10, which gives the macroscopic description of $(R,D)$-thin non-compact spaces.
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space (see Definition 4.2). Let $r$ be a ray in $X$ and let $y\in X$ with $\mathsf{d}(r,y)>D$. Let $t>0$ and let $\gamma$ be a curve connecting $y$ and $r(t)$. Then, for every $s\in[0,t]$, it holds $\mathsf{d}(r(s),\gamma)\leq 2R$.
@@
We parametrize $\gamma$ by arc length in such a way that $\gamma(0)=r(t)$ and $\gamma(\mathrm{L}(\gamma))=y$. Since $X$ is $(R,D)$-thin and $\mathsf{d}(y,r)>D$, it follows that $\pi_{r}(\gamma(\mathrm{L}(\gamma)))\in[r(0),r(R)]$. By Lemma 4.5, $\pi_{r}(\gamma)$ contains a $3D$-net of $[r(R),r(t)]$. Using again that $X$ is $(R,D)$-thin, it follows that $[r(R),r(t)]$ is in a $7D$-neighbourhood of $\gamma$. The statement then follows. ∎
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space (see Definition 4.2). If there are two divergent rays in $X$, then $X$ contains a line.
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space (see Definition 4.2). Let $r$ be a ray in $X$. Let $y\in X$ be a point such that $\mathsf{d}(y,r)>D$. Let $r_{y}$ be a ray obtained as a limit of the segments connecting $y$ and $r(t_{j})$, for some sequence $t_{j}\to+\infty$. Then, $r$ is contained in the $2R$-neighbourhood of $r_{y}$.
@@
The next proposition shows that an $(R,D)$-thin non-compact metric space is contained in a neighbourhood of controlled width of a ray or a line. We recall that a similar result is proved in [TrianglesKapo, Section 5]. The main difference is that Proposition 4.10 below also gives an explicit bound on the width of the aforementioned neighbourhood.
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space (see Definition 4.2). Assume that $X$ is non-compact and it does not contain a line. There exists a ray $r$ in $X$ whose $10R$-neighbourhood contains $X$.
@@
We now prove the statement of the lemma. If $X$ is contained in the $10R$-neighbourhood of $r$, there is nothing to prove. Otherwise, we claim that there exists an element $y\in X$ at maximal distance from $r$. Let $y_{i}\in X$ be a sequence maximizing the distance from $r$. We can assume that $9R\leq\mathsf{d}(y_{i},r)\leq D^{\prime}(X)$. Since $X$ is $(R,D)$-thin, all projections of $y_{i}$ on $r$ are contained in $[r(0),r(R)]$. It follows that the sequence $\{y_{i}\}_{i\in\mathbb{N}}$ is precompact, so that there exists a limit point $y\in X$ at maximal distance from $r$.
@@
We now turn our attention to the study of compact $(R,D)$-thin metric spaces. The key result in this sense is Proposition 4.15. We first need a few lemmas.
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $\bar{l}$ be a segment starting from $l(0)$ with $\pi_{l}(\bar{l}(\mathrm{L}(\bar{l})))\subset[l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$. If $\bar{l}$ has a projection on $l$ lying in $[l(3R),l(\mathrm{L}(l)-3R)]$, then $\pi_{l}(\bar{l})$ contains a $3D$-net of $[l(R),l(\mathrm{L}(l)-R)]$.
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $p\in X$ be a point with $\mathsf{d}(p,l)\geq 10R$ and let $\bar{l}$ be a segment connecting $p$ to $l(0)$. If $\bar{l}$ has a projection on $l$ lying in $[l(3R),l(\mathrm{L}(l)-3R)]$, then $\mathrm{L}(\bar{l})>\mathrm{L}(l)$.
@@
Since $\mathsf{d}(p,l)\geq 10R$ and $X$ is $(R,D)$-thin, it follows that $\pi_{l}(p)\subset[l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$. By Lemma 4.11, there is a point $\bar{l}(s_{0})$ which is $7D$-close to $l(\mathrm{L}(l)-R)$. Using triangle inequality, this implies
@@
Let $(X,\mathsf{d})$ be a compact $(R,D)$-thin metric space. Let $l$ be a segment of maximal length in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $p\in X$ be a point with $\mathsf{d}(p,l)\geq 10R$ and let $\gamma_{1},\gamma_{2}$ be segments connecting $p$ respectively to $l(\mathrm{L}(l))$ and $l(0)$. Let $b_{1},b_{2}\in\{\gamma_{1},\gamma_{2},l\}$ with $b_{1}\neq b_{2}$. Then $\pi_{b_{1}}(b_{2})\subset[b_{1}(0),b_{1}(3R)]\cup[b_{1}(\mathrm{L}(b_{1})-3R),b_{1}(\mathrm{L}(b_{1})]$.
@@
We now consider the case $b_{1}\in\{\gamma_{1},\gamma_{2}\}$ and $b_{2}=l$. If the statement fails, $l$ has a projection on $b_{1}$ lying in $(b_{1}(3R),b_{1}(\mathrm{L}(b_{1})-3R))$. Consider the orientation of $l$ such that $l(0)\in b_{1}$. Then, $\pi_{b_{1}}(l(\mathrm{L}(l)))\subset[b_{1}(0),b_{1}(R)]\cup[b_{1}(\mathrm{L}(b_{1})-R),b_{1}(\mathrm{L}(b_{1})]$ since otherwise, using that $X$ is $(R,D)$-thin, it follows that $l$ is shorter than $b_{1}$. Hence, by Lemma 4.11, there exists $s>0$ such that $l(s)$ is $R+7D$-close to $p$. This contradicts that $\mathsf{d}(p,l)\geq 10R$.